OpenStudy (anonymous):

What is the x-coordinate of the solution to the system of equations? 3x – 2y = –30 6x + 3y = 3

OpenStudy (kidrah69):

\(\huge\color{orange}{Welcome~To~OpenStudy}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions

OpenStudy (kidrah69):

Elimination will be the easiet here :) Do you know how to use the elimination method?

OpenStudy (triciaal):

Using substitution method from equation 2 2x + y = 1 so y = 1-2x substitute for y in equation 1 3x - 2(1-2x) = -30 7x = -28

OpenStudy (anonymous):

thanks

OpenStudy (triciaal):

you are welcome

OpenStudy (anonymous):

could u helpme with more?